# Cooperative Sheaf Neural Networks

### v1

@@ 4 Cooperative Sheaf Neural Networks
In this section, we leverage the sheaf Laplacians in Definition 3.2 to introduce Cooperative Sheaf Neural Networks (CSNNs). To exploit the asymmetric communication induced by directed sheaves, we first convert our input (undirected) graph into a directed one in the following way: every time there is an edge between nodes $i$ and $j$, an arbitrary direction is chosen to form the graph $G$, then we also consider the reversed graph $\overline{G}$. Finally, we use a sheaf $(G\cup\overline{G},{\mathcal{F}})$ whose restriction maps $\mathcal{F}_{i\unlhd ij}$ and $\mathcal{F}_{i\unlhd ji}$ control the information from $i$ and to $i$, respectively.
Our goal is to create a sheaf diffusion model capable of exhibiting cooperative behavior. In other words, nodes should be able to decide how they participate in message passing, choosing whether to broadcast their information or to listen from the neighbors. For undirected sheaf models, it is easy to see that this behavior is unachievable since both the message from $i$ to $j$ through an undirected edge $e$ and its converse depend directly on $\mathcal{F}_{i\unlhd e}^{\top}$ — and setting $\mathcal{F}_{i\unlhd e}^{\top}$ to zero cancels out messages in both directions (c.f., Equation 1).
We design CSNN to achieve this flexible behavior by composing the out-degree and the transposed in-degree sheaf Laplacians. For each node $i$, we will define a source conformal map ${\mathbf{S}}_{i}$ and a target conformal map ${\mathbf{T}}_{i}$, for all neighbor $j$ of $i$. In this case, out-degree sheaf Laplacian simplifies to
$L^{\text{out}}_{\mathcal{F}}(\mathbf{X})_{i}$ $:=\textstyle\sum_{j\in N(i)}\left({\mathbf{S}}_{i}^{\top}{\mathbf{S}}_{i}\mathbf{x}_{i}-{\mathbf{T}}_{i}^{\top}{\mathbf{S}}_{j}\textbf{x}_{j}\right),$ (4)
@@
Note these matrices have a block structure, with diagonals $({L^{\text{in}}_{\mathcal{F}}})^{\top}_{ii}=\sum{\mathbf{T}}_{i}^{\top}{\mathbf{T}}_{i}$, $({L^{\text{out}}_{\mathcal{F}}})_{ii}=\sum{\mathbf{S}}_{i}^{\top}{\mathbf{S}}_{i}$, and remaining blocks $(L^{\text{in}}_{{\mathcal{F}}})^{\top}_{ij}=(L^{\text{out}}_{{\mathcal{F}}})_{ij}=-{\mathbf{T}}^{\top}_{i}{\mathbf{S}}_{j}$. In practice, we use their normalized versions
@@
where $D_{\text{in}},D_{\text{out}}$ are the block-diagonals of the in and out Laplacians, respectively. We point out that, since conformal maps are of the form ${\mathbf{S}}_{i}=C_{{\mathbf{S}}_{i}}Q_{i}$ and ${\mathbf{T}}_{i}=C_{{\mathbf{T}}_{i}}R_{i}$, for some orthogonal matrices $Q_{i}$, $R_{i}$ and scalars $C_{{\mathbf{S}}_{i}},\>C_{{\mathbf{T}}_{i}}$, the block diagonals are just scalars times an identity matrix, hence the normalization is both numerically stable and computationally efficient.
The CSNN model is an adaptation of the NSD model from Bodnar et al. (2022), which is given by
$\mathbf{X}_{t+1}=(1+\varepsilon)\mathbf{X}_{t}-\sigma(\Delta_{{\mathcal{F}}(t)}(\mathbf{I}\otimes\mathbf{W}_{1,t})\mathbf{X}_{t}\mathbf{W}_{2,t}),$
where $\Delta_{{\mathcal{F}}(t)}$ is the sheaf Laplacian Equation 1 for undirected graphs. Concisely, we define a CSNN layer from the NSD model substituting the sheaf Laplacian of the undirected graph by the composition of normalized in- and out-Laplacians:
$\textbf{X}_{t+1}=(1+\varepsilon)\textbf{X}_{t}-\sigma((\Delta_{{\mathcal{F}}(t)}^{in})^{\top}\Delta_{{\mathcal{F}}(t)}^{out}(\textbf{I}\otimes\textbf{W}_{1,t})\textbf{X}_{t}\textbf{W}_{2,t}),$ (6)
where $\varepsilon\in[-1,1]^{nd}$, $\sigma$ is an activation, and ${\mathbf{W}}_{1}\in{\mathbb{R}}^{d\times d},{\mathbf{W}}_{2}\in{\mathbb{R}}^{h\times h}$ are weight matrices. We provide a pseudo-code on how we implement the CSNN model in Algorithm 1.
Algorithm 1 Cooperative Sheaf Diffusion
Input: Node features $X_{0}\in{\mathbb{R}}^{n\times s}$, graph $G$, number of layers $L$.
Output: Class probabilities for each node $X_{\mathrm{final}}$
1:$\mathbf{X}_{0}=\sigma(X_{0}W_{0}+b_{0})$triangleright $\mathbf{X}_{0}\in{\mathbb{R}}^{nd\times h}$
2:for $\ell=0,\dots,L-1$ do
3:     ${\mathbf{S}}_{i}=\eta(\mathbf{X}_{\ell},G,i),\>\forall i\in V$triangleright Learn source maps
4:     ${\mathbf{T}}_{i}=\phi(\mathbf{X}_{\ell},G,i),\>\forall i\in V$triangleright Learn target maps
5:     $L_{{\mathcal{F}}}^{\text{in}},L_{{\mathcal{F}}}^{\text{out}}=\mathrm{LaplacianBuilder}(\{{\mathbf{S}}_{i}\}_{i=1}^{n},\{{\mathbf{T}}_{i}\}_{i=1}^{n})$
6:     $\mathbf{X}_{\ell+1}=\mathrm{Diffusion}(\mathbf{X}_{\ell},L_{{\mathcal{F}}}^{\text{in}},L_{{\mathcal{F}}}^{\text{out}})$triangleright Equation 6
7:end for
8:Return: $\mathrm{READOUT}(\mathbf{X}_{L})$
@@
The first step of each layer $\ell$ consists in computing the conformal maps ${\mathbf{T}}_{i,\ell}$ and ${\mathbf{S}}_{i,\ell}$. We do this through learnable functions ${\mathbf{S}}_{i,\ell}=\eta(G,\mathbf{X}_{\ell},i)$ and ${\mathbf{T}}_{i,\ell}=\phi(G,\mathbf{X}_{\ell},i)$, $\forall i\in V$. As in Bamberger et al. (2024), we can use any neural network to learn the restriction maps. In detail, we use Householder reflections (Mhammedi et al., 2017; Obukhov, 2021) to compute orthogonal maps and multiply them by a learned positive constant for each node. Finally, the Laplacians $(\Delta_{{\mathcal{F}}}^{in})^{\top}$ and $\Delta_{{\mathcal{F}}}^{out}$ are computed. In light of Proposition 4.1, it is clear that this modification was intended to create a model in which a node can ignore its neighbors or choose not to send them information.
In particular, CSNN allows each node $i$ to play one of the four following roles in information diffusion:
  • STANDARD: This is the standard message-passing paradigm. The node $i$ propagates information from and to its neighbors. Mathematically, this corresponds to ${\mathbf{S}}_{i},{\mathbf{T}}_{i}\neq 0$.
  • LISTEN: The node $i$ receives information from broadcasting neighbors, i.e. ${\mathbf{S}}_{i}=0$ and ${\mathbf{T}}_{i}\neq 0$.
  • BROADCAST: Node $i$ propagates information to the neighbors that listen, i.e. ${\mathbf{T}}_{i}=0$ and ${\mathbf{S}}_{i}\neq 0$.
  • ISOLATE: The node $i$ neither listens nor broadcasts, i.e. ${\mathbf{S}}_{i}=0={\mathbf{T}}_{i}$.
@@
Figure 2: ${\mathbf{O}}_{i}=0$ creates the effect of isolating the node $i$. Directed edges provide the possibility of considering ${\mathbf{S}}_{i}=0$ and ${\mathbf{T}}_{i}=0$, performing listen and broadcast, separately.
@@
We highlight the importance of considering the directions: for undirected graphs, we have a single map ${\mathbf{O}}_{i}$ for each node $i$, where ${\mathbf{O}}_{i}=0$ could only mean that $i$ does not communicate at all. In other words, the possible actions are only standard and isolate. We illustrate this in Figure 2.
Goal. Ideally, our model must satisfy the following: if $i$ does not listen, then its update cannot depend of $\mathbf{x}_{j}$, $\forall j\in V,j\neq i$. If $i$ has a non-broadcasting neighbor $k$, then its update cannot depend on $\mathbf{x}_{k}$. The composition of out- and (transposed) in-Laplacians has this property, as we prove in Proposition 4.1.

### v2

@@ 4 Cooperative Sheaf Neural Networks
In this section, we leverage the sheaf Laplacians in Definition˜3.3 to propose Cooperative Sheaf Neural Networks (CSNNs), an SNN which allows nodes to independently to decide how they participate in message diffusion, choosing whether to broadcast their information and/or to listen from the neighbors. To exploit the asymmetric communication induced by sheaves over directed graphs, we convert our input undirected graph into a directed one by replacing undirected edges with a pair of directed ones.
We design CSNN’s diffusion mechanism by composing the out-degree and the transposed in-degree sheaf Laplacians. In practice, we use their normalized versions
where $D_{\text{in}},D_{\text{out}}$ are the block-diagonals of the in and out Laplacians, respectively.
We define a CSNN layer by augmenting the Euler discretization of our novel heat equation $\dot{\mathbf{X}}=(\Delta_{{\mathcal{F}}}^{in})^{\top}\Delta_{{\mathcal{F}}}^{out}\mathbf{X}$ with linear transformations and a nonlinear activation function $\sigma$: $\mathbf{X}_{t+1}=(1+(\bm{1}_{n\times h}\otimes\varepsilon))\odot\mathbf{X}_{t}-\sigma((\Delta_{{\mathcal{F}}(t)}^{in})^{\top}\Delta_{{\mathcal{F}}(t)}^{out}(\mathbf{I}_{n}\otimes\mathbf{W}_{1,t})\mathbf{X}_{t}\mathbf{W}_{2,t}),$ (6) where $\bm{1}_{n\times h}$ is an $n$-by-$h$ matrix of ones, $\varepsilon\in[-1,1]^{d\times 1}$, and ${\mathbf{W}}_{1,t}\in{\mathbb{R}}^{d\times d}$ and ${\mathbf{W}}_{2,t}\in{\mathbb{R}}^{h\times h}$ are learned matrices responsible for mixing node features and channels, respectively.
Efficient implementation.
For computational efficiency, we use flat vector bundles to define both the in- and out- degree sheaf Laplacians. More precisely, for each node $i$, we define a source conformal map ${\mathbf{S}}_{i}$ and a target conformal map ${\mathbf{T}}_{i}$ for all neighbor $j$ of $i$. Thus, out-degree sheaf Laplacian simplifies to
$L^{\text{out}}_{\mathcal{F}}(\mathbf{X})_{i}$ $:=\textstyle\sum_{j\in N(i)}\left({\mathbf{S}}_{i}^{\top}{\mathbf{S}}_{i}\mathbf{x}_{i}-{\mathbf{T}}_{i}^{\top}{\mathbf{S}}_{j}\mathbf{x}_{j}\right),$ (7)
@@
Note these matrices have a block structure, with diagonals $({L^{\text{in}}_{\mathcal{F}}})^{\top}_{ii}=\sum{\mathbf{T}}_{i}^{\top}{\mathbf{T}}_{i}$, $({L^{\text{out}}_{\mathcal{F}}})_{ii}=\sum{\mathbf{S}}_{i}^{\top}{\mathbf{S}}_{i}$, and remaining blocks $(L^{\text{in}}_{{\mathcal{F}}})^{\top}_{ij}=(L^{\text{out}}_{{\mathcal{F}}})_{ij}=-{\mathbf{T}}^{\top}_{i}{\mathbf{S}}_{j}$. We point out that, since conformal maps are of the form ${\mathbf{S}}_{i}=C_{{\mathbf{S}}_{i}}Q_{i}$ and ${\mathbf{T}}_{i}=C_{{\mathbf{T}}_{i}}R_{i}$, for some orthogonal matrices $Q_{i}$, $R_{i}$ and scalars $C_{{\mathbf{S}}_{i}},\>C_{{\mathbf{T}}_{i}}$, the block diagonals are just scalars times an identity matrix, hence the normalization is both numerically stable and computationally efficient.
@@
The first step of each layer $t$ consists of computing the conformal maps ${\mathbf{T}}_{i,t}$ and ${\mathbf{S}}_{i,t}$. We do this through learnable functions ${\mathbf{S}}_{i,t}=\eta(G,\mathbf{X}_{t},i)$ and ${\mathbf{T}}_{i,t}=\phi(G,\mathbf{X}_{t},i)$, $\forall i\in V$. As in prior works on SNNs Bodnar et al. (2022); Bamberger et al. (2025), we use neural networks to learn the restriction maps. In addition, we use Householder reflections (Mhammedi et al., 2017; Obukhov, 2021) to compute orthogonal maps and multiply them by a learned positive constant for each node.
4.1 Analysis
@@
Figure 2: ${\mathbf{O}}_{i}=0$ creates the effect of isolating the node $i$. Directed edges provide the possibility of performing LISTEN and PROPAGATE, separately.
We now show that CSNN can achieve cooperative behavior, characterize its receptive field, and prove that appropriate conformal maps can help CSNNs handle long-range interactions.
We note that CSNN allows each node $i$ drift from the STANDARD behavior (${\mathbf{S}}_{i},{\mathbf{T}}_{i}\neq 0$) by zeroing-out its conformal maps. Setting ${\mathbf{T}}_{i}\neq 0$ drives $i$ to LISTEN. Setting ${\mathbf{S}}_{i}\neq 0$ corresponds to PROPAGATE. Finally, ${\mathbf{S}}_{i}={\mathbf{T}}_{i}=0$ implies ISOLATE.
@@
We highlight the importance of considering the directions: for undirected graphs, there is a single map ${\mathbf{O}}_{i}$ for each node $i$, where ${\mathbf{O}}_{i}=0$ could only mean that $i$ does not communicate at all. In other words, the possible actions are only STANDARD and ISOLATE. We illustrate this in Figure˜2.
Thus, to show a model achieves cooperative behavior, we must ensure the following: (a) if $i$ does not listen, then its update cannot depend on $\mathbf{x}_{j}$, $\forall j\in V,j\neq i$; and (b) if $i$ has a non-propagating neighbor $k$, then its update cannot depend on $\mathbf{x}_{k}$. Proposition˜4.1 shows CSNN satisfies these conditions.